# Fundamentals and advances in transverse thermoelectrics

### v2

@@ II.1 Transverse thermoelectric effects with time-reversal symmetry breaking
where $\mu_{0}$ is the vacuum permeability and $Q_{H}$ ($Q_{M}$) is the proportionality factor of the $H$-dependent ($M$-dependent) term. Equation (1) shows that the anomalous Nernst coefficient $S_{\rm ANE}$ ($=Q_{M}\mu_{0}M_{\rm s}$ with $M_{\rm s}$ being the saturation magnetization) can be extracted by extrapolating the $H$ dependence of $S_{\rm T}$ from the high field region, in which $M$ is saturated, to zero field. Thus, if ${\bf M}$ is aligned in one direction, ANE can generate a thermoelectric voltage and power even without applying an external magnetic field. Note that the magnitude of the transverse thermopower due to ANE is determined by $S_{\rm ANE}$, not $Q_{M}$; $Q_{M}$ is just a phenomenological parameter defined for convenience to separate ANE from ONE. Based on Eq. (1), $S_{\rm ANE}$ is often compared in terms of $M$ to discuss the scaling behavior. However, since ANE and related transport properties are estimated for uniformly magnetized materials, they should be compared in terms of $M_{\rm s}$, not $M$. In fact, $S_{\rm ANE}$ is not correlated with $M_{\rm s}$ in many magnetic materials; the scaling behavior does not hold even for simple ferromagnetic metals, such as Fe, Ni, and Co [37].
@@
[…]
@@ III.1 Revising performance of longitudinal thermoelectric generator
$\begin{pmatrix}q_{x}\\ q_{y}\end{pmatrix}$ $=$ $T{S}_{\nu}\begin{pmatrix}j_{x}\\ j_{y}\end{pmatrix}-{\kappa}_{\nu}\begin{pmatrix}\nabla_{x}T\\ \nabla_{y}T\end{pmatrix},\qquad$ (4)
@@
^$\widehat{\cal B}=\begin{pmatrix}-TS/R&K^{\prime}\\ -1/R&S/R\end{pmatrix}.$
@@ III.2 Performance of transverse thermoelectric generator
$\begin{pmatrix}q_{x}\\ q_{y}\end{pmatrix}$ $=$ $T\widehat{S}\begin{pmatrix}j_{x}\\ j_{y}\end{pmatrix}-\widehat{\kappa}\begin{pmatrix}\nabla_{x}T\\ \nabla_{y}T\end{pmatrix},\qquad$ (23)
@@
[…]
@@
Figure 5: Relations between $z_{xy}^{\rm(i)}$, $z_{xy}^{\rm(a)}$, and $\overline{\eta}$ for various values of $zT$. $z_{xy}^{\rm(i)}$ ($z_{xy}^{\rm(a)}$) and $zT$ denote the transverse figure of merit in the isothermal (adiabatic) condition and the longitudinal figure of merit for the Seebeck effect, respectively.
@@ III.3 Performance of transverse thermoelectric refrigerator
We first need to obtain the differential equation that determines temperature distribution in the device. The rate of heat evolution is given by $-{\bm{\nabla}}\cdot{\bf q}_{E}$, where ${\bf q}_{E}={\bf q}+\phi{\bf j}$ is the energy current density with $\phi$ being the electrochemical potential. Then, assuming the steady-state condition, the temperature distribution in the device is determined by [140]
$-{\bm{\nabla}}\cdot{\bf q}_{E}=-{\bm{\nabla}}\cdot{\bf q}+{\bf E\cdot j}=0.$ (67)
@@
[…]
@@ III.4 Performance of transverse active cooler
where $a=2S_{yx}\langle j_{y}\rangle_{\rm s}/\kappa_{xx}$ and $b=\rho_{yy}\langle j_{y}\rangle_{\rm s}^{2}/\kappa_{xx}$. Note that this expression is also obtained by interchanging $T_{\rm h}$ and $T_{\rm c}$ in Eq. (74). Then, substituting this temperature distribution into Eq. (68), we obtain the rate of heat removal from the hot reservoir ${\cal Q}_{\rm h}=l_{y}l_{z}q_{x}(x=0)$ as ${\cal Q}_{\rm h}=K\Delta T-({R}/{2})J_{y}^{2}-T_{\rm c}S_{yx}(l_{y}/l_{x})J_{y}$, where $J_{y}=l_{x}l_{z}\langle j_{y}\rangle_{\rm s}$, and $K$ and $R$ are defined below Eq. (47). Completing the square with respect to $J_{y}$ in the above equation, we obtain ${\cal Q}_{\rm h}=K\Delta T-({R}/{2})\big{(}J_{y}-J_{y}^{*}\big{)}^{2}+({T_{\rm c}^{2}S_{yx}^{2}}/{2R})(l_{y}/l_{x})^{2}$, where $J_{y}^{*}=-T_{\rm c}S_{yx}(l_{y}/l_{x})^{2}/2R$. Then, by setting $J_{y}=J_{y}^{*}$, the maximum heat removal from the hot reservoir ${\cal Q}_{\rm h}^{*}$ is obtained as ${\cal Q}_{\rm h}^{*}=K\Delta T+({T_{\rm c}^{2}S_{yx}^{2}}/{2R})(l_{y}/l_{x})^{2}$. Now we introduce the heat current density $q_{\rm h}^{*}={\cal Q}^{*}_{\rm h}/(l_{y}l_{z})$, and rewrite the above equation as $q_{\rm h}^{*}=\kappa_{xx}\left(({\Delta T}/{l_{x}})\right)+({T_{\rm c}^{2}S_{yx}^{2}})/({2\rho_{yy}l_{x}})$. Then, defining ~$\widetilde{\nabla}_{x}T=\Delta T/l_{x}$ which is positive by definition, the above equation can be represented as
@@ IV.4 Transverse active cooler
Finally, we discuss the performance of the active cooler based on the off-diagonal Peltier effect. Let us first consider the case of isothermal condition in the transverse direction. In this case, as in the previous section, we need to determine the temperature distribution by solving Eq. (72). Then, we obtain the temperature distribution expressed by Eq. (82), where $a=0$ and $b=\rho_{yy}j_{y}^{2}/\kappa_{xx}$ should be noted. Then, substituting this temperature distribution into Eq. (68), we obtain the rate of heat removal from the hot reservoir ${\cal Q}_{\rm h}=l_{y}l_{z}q_{x}(x=0)$ as ${\cal Q}_{\rm h}=K\Delta T-({R}/{2})J_{y}^{2}+T_{\rm h}S_{yx}(l_{y}/l_{x})J_{y}$, where $J_{y}=l_{x}l_{z}j_{y}$. Completing the square with respect to $J_{y}$ in the above equation, we obtain ${\cal Q}_{\rm h}=K\Delta T-({R}/{2})\big{(}J_{y}-J_{y}^{*}\big{)}^{2}+({T_{\rm h}^{2}S_{yx}^{2}}/{2R})(l_{y}/l_{x})^{2}$, where $J_{y}^{*}=-{T_{\rm h}S_{yx}}(l_{y}/l_{x})/2R$. Then, by setting $J_{y}=J_{y}^{*}$, the maximum heat removal from the hot reservoir ${\cal Q}_{\rm h}^{*}$ is obtained as ${\cal Q}_{\rm h}^{*}=K\Delta T+({T_{\rm h}^{2}S_{yx}^{2}}/{2R})(l_{y}/l_{x})^{2}$. Now we define $q_{\rm h}^{*}={\cal Q}^{*}_{\rm h}/(l_{y}l_{z})$, and rewrite the above equation as $q_{\rm h}^{*}=\kappa_{xx}\left({\Delta T}/{l_{x}}\right)+({T_{\rm h}^{2}S_{yx}^{2}}/{2\rho_{yy}l_{x}})$. Then, with the definition of ~$\widetilde{\nabla}_{x}T=\Delta T/l_{x}$, the above equation can be transformed to Eq. (83), in which the effective thermal conductivity for the present case is given by
@@ V Measurement of transverse thermopower
[…]
@@ V.1 Measurement configuration for Nernst effect in thin films
[…]
@@ V.2 Effect of thermal boundary conditions
We now experimentally demonstrate the influence of sample dimensions on the transverse thermopower using a bulk polycrystalline SmCo5 slab. The SmCo5 block, commercially available from Magfine Corporation, Japan, was cut into two rectangular slabs with dimensions along the $x$, $y$, and $z$ directions respectively being 10.0, 7.7, and 0.8 mm and 10.0, 2.1, and 0.8 mm using a diamond wire saw, where only the length along the $y$ direction, $d$, is significantly different. The slabs have an easy axis of magnetization along $z$ direction. The measurement configuration is illustrated in Fig. 7(c). The two SmCo5 slabs were bridged between two heat baths separated by 8.0 mm and firmly attached to the heat baths. To apply a temperature difference between the heat baths, chip heaters were attached to the hot side. The sample holder was fixed at the center of a superconducting magnet, where a magnetic field ranging from $\mu_{0}H=5~{\rm T}$ to $-5~{\rm T}$ was applied along the $z$ direction. The $H$ dependence of the transverse voltage $V_{y}$ along the $y$ direction was recorded by attaching Cu wires at the ends of the slabs at various values of the temperature difference. To accurately measure the temperature gradient on the SmCo5 slabs, the top surface of the slabs was coated with insulating black ink with an emissivity of >0.94>0.94> 0.94 and steady-state temperature images were captured using an infrared camera.
@@
Figure 7(d) shows the $H$ dependence of the transverse electric field $V_{y}/d$ for the SmCo5 slabs, recorded under a longitudinal temperature gradient of $|\nabla_{x}T|\sim 1\times 10^{3}~{\rm K/m}$. In both the slabs, the transverse electric field changed its sign with large hysteresis with respect to the $H$ reversal. The $H$ dependence of $V_{y}/d$ corresponds to the magnetization curve of the SmCo5 permanent magnet, indicating the appearance of ANE. To determine the magnitude of the $M$-dependent transverse thermopower for the SmCo5 slabs, the $M$-odd-dependent component of $V_{y}/d$ under the saturation state, $V_{\rm odd}/d$, was extracted for each $|\nabla_{x}T|$ value by extrapolating the signal in the high $H$ region to zero field. Here, the finite $H$-linear component was observed only for the slab with $d=2.1~{\rm mm}$, suggesting the presence of the Seebeck voltage between the SmCo5/Cu-wire contacts induced by the ordinary ($H$-dependent) THE in SmCo5 [Fig. 7(d)]. Although the SmCo5 slabs were cut from the same ingot, they exhibited different $V_{\rm odd}/d$ values [Fig. 7(e)]. The $M$-dependent transverse thermopower for the SmCo5 slab with $d=2.1~{\rm mm}$ ($d=7.7~{\rm mm}$) was found to be $5.1~\mu{\rm V/K}$ ($4.1~\mu{\rm V/K}$); the difference in the transverse thermopower is much larger than measurement errors. Recall the fact that the slab with $d=2.1~{\rm mm}$ ($d=7.7~{\rm mm}$) is closer to the adiabatic (isothermal) condition. Therefore, the transverse thermopower under the adiabatic condition is expected to be overestimated compared to that under the isothermal condition because of the same sign of $S_{xy}$ and the correction term induced by the Seebeck effect and anomalous ($M$-dependent) THE. This expectation is consistent with the observed difference in the $M$-dependent transverse thermopower. Here, we note that the correction term discussed theoretically in Sec. III is determined by the absolute Seebeck coefficient of a magnetic material (i.e. SmCo5), while the observed difference in the transverse thermopower in Fig. 7 depends on the relative Seebeck coefficient between SmCo5 and Cu. However, since the Seebeck coefficient of Cu ($<2~{\rm\mu V/K}$) is much smaller than that of SmCo5 ($-19~{\rm\mu V/K}$ [63]) at room temperature, the magnitude and sign of the correction term in the present setup is determined mainly by the transport properties of SmCo5. The experimental results shown here offer guidelines for determining the appropriate sample aspect ratio and measurement setup for transverse thermopower measurements, minimizing parasitic signals from the Seebeck effect and THE. We also note that considering typical sample dimensions, the estimation of $S_{\rm ANE}$ from the measurement of the anomalous Ettingshausen effect and the Onsager reciprocal relation reduces the influence of the correction term [29].

### v3

@@ II.1 Transverse thermoelectric effects with time-reversal symmetry breaking
where $\mu_{0}$ is the vacuum permeability and $Q_{H}$ ($Q_{M}$) is the proportionality factor of the $H$-dependent ($M$-dependent) term. Equation (1) shows that the anomalous Nernst coefficient $S_{\rm ANE}$ ($Q_{M}\mu_{0}M_{\rm s}$ with $M_{\rm s}$ being the saturation magnetization) can be extracted by extrapolating the $H$ dependence of $S_{\rm T}$ from the high field region, in which $M$ is saturated, to zero field. Thus, if $\bf M$} is aligned in one direction, ANE can generate a thermoelectric voltage and power even without applying an external magnetic field. Note that the magnitude of the transverse thermopower due to ANE is determined by $S_{\rm ANE}$, not $Q_{M}$; $Q_{M}$ is just a phenomenological parameter defined for convenience to separate ANE from ONE. Based on Eq. (1), $S_{\rm ANE}$ is often compared in terms of $M$ to discuss the scaling behavior. However, since ANE and related transport properties are estimated for uniformly magnetized materials, they should be compared in terms of $M_{\rm s}$, not $M$. In fact, $S_{\rm ANE}$ is not correlated with $M_{\rm s}$ in many magnetic materials; the scaling behavior does not hold even for simple ferromagnetic metals, such as Fe, Ni, and Co [37].
@@
where $\rho_{\rm M(TE)}$ and $S_{\rm M(TE)}$ are the longitudinal resistivity and Seebeck coefficient of the magnetic (thermoelectric) material, respectively, and $r$ is the size ratio determined by the dimensions of the magnetic and thermoelectric materials [71]. Here, we consider a situation in which magnetic and thermoelectric materials are separated by an insulating layer to avoid the shunting effect. When the in-plane areas of the magnetic and thermoelectric materials are the same, $r$ is determined by the thickness ratio of the two: $r=t_{\rm TE}/t_{\rm M}$ with $t_{\rm M(TE)}$ being the thickness of the magnetic (thermoelectric) material. The second term on the right-hand side of Eq. (3) represents the contribution of SAHE. By optimizing the combination and size ratio of the magnetic and thermoelectric materials, it is possible to obtain a much larger transverse thermopower than $S_{\rm ANE}$ alone. When a thin film of a magnetic material is combined with a slab of a thermoelectric material, $r$ becomes very large, and $S_{\rm T}$ approaching $100~\mu{\rm V/K}$ has been observed [71]. When a slab of a magnetic material is combined with a slab of a thermoelectric material, $S_{\rm T}$ is limited to around 10-$20~\mu{\rm V/K}$ due to smaller $r$ but such all-bulk hybrid structure is more suitable for increasing output power [72]. Recently, direct-contact SAHE has also been demonstrated, in which a magnetic material is directly deposited on the surface of a thermoelectric material, rather than forming a closed circuit [73]. Direct-contact SAHE functions with a simpler device structure, but its transverse thermopower decreases due to the shunting effect compared to the value for the structure depicted in Fig. 3(d). Since there is a vast amount of research data accumulated over many years on AHE and the Seebeck effect, there is still plenty of room for improving the performance of SAHE. However, its versatility is lower than that of ANE; whereas ANE works in both in-plane and perpendicularly magnetized configurations, SAHE works only in the perpendicularly magnetized configuration. Furthermore, even if $S_{\rm T}$ is enhanced, $z_{xy}T$ and power generation efficiency of SAHE are not always higher than those of ANE [74].
@@ III.1 Revising performance of longitudinal thermoelectric generator
$\begin{pmatrix}q_{x}\\ q_{y}\end{pmatrix}$ $=$ $T{S}_{\nu}\begin{pmatrix}j_{x}\\ j_{y}\end{pmatrix}-{\kappa}_{\nu}\begin{pmatrix}\nabla_{x}T\\ \nabla_{y}T\end{pmatrix}$$\qquad$ (4)
@@
$\widehat{\cal B}=\begin{pmatrix}-TS/R&K^{\prime}\\ -1/R&S/R\end{pmatrix}$.
@@ III.2 Performance of transverse thermoelectric generator
$\begin{pmatrix}q_{x}\\ q_{y}\end{pmatrix}$ $=$ $T\widehat{S}\begin{pmatrix}j_{x}\\ j_{y}\end{pmatrix}-\widehat{\kappa}\begin{pmatrix}\nabla_{x}T\\ \nabla_{y}T\end{pmatrix}$$\qquad$ (23)
@@
${\rho}_{yy}^{\rm(a)}$ $=$ $\rho_{yy}+\frac{T{S}_{yy}^{\rm(a)\,2}}{{\kappa}^{\rm(a)}_{yy}}+\frac{T{S}^{\rm(a)}_{xy}{S}^{\rm(a)}_{yx}}{{\kappa}^{\rm(a)}_{xx}}$ (36)
$=$ $\rho_{yy}+\frac{T{S}_{yy}^{\rm(a)\,2}}{{\kappa}^{\rm(a)}_{yy}}-\frac{T{S}_{yx}^{\rm(a)\,2}}{{\kappa}^{\rm(a)}_{xx}}$
@@
Figure 5: Relations between $z_{xy}^{\rm(i)}$, $z_{xy}^{\rm(a)}$, and $\overline{\eta}$ for various values of $zT$. $z_{xy}^{\rm(i)}$ ($z_{xy}^{\rm(a)}$) and zTzT denote the transverse figure of merit in the isothermal (adiabatic) condition and the longitudinal figure of merit for the Seebeck effect, respectively.
@@ III.3 Performance of transverse thermoelectric refrigerator
We first need to obtain the differential equation that determines temperature distribution in the device. The rate of heat evolution is given by ${\bm{\nabla}}\cdot{\bf q}_{E}$, where ${\bf q}_{E}={\bf q}+\phi{\bf j}$ is the energy current density with $\phi$ being the electrochemical potential. Then, assuming the steady-state condition, the temperature distribution in the device is determined by [140]
${\bm{\nabla}}\cdot{\bf q}_{E}=-{\bm{\nabla}}\cdot{\bf q}+{\bf E\cdot j}=0.$ (67)
@@
Since the temperature distribution is obtained in the same way, we now proceed exactly in the same manner as in the preceding subsection, and obtain the rate of the heat removal from the cold reservoir as ${\cal Q}_{\rm c}=-{K}^{\rm(a)}\Delta T-({R^{\rm(b)}}/{2})J_{y}^{2}-T_{\rm h}{S}^{\rm(a)}_{yx}(l_{y}/l_{x})J_{y}$, where ${K}^{\rm(a)}={\kappa}^{\rm(a)}_{xx}(l_{y}l_{z}/l_{x})$, $R^{\rm(b)}=\rho^{\rm(b)}_{yy}l_{y}/(l_{x}l_{z})$, and $J_{y}=l_{x}l_{z}\langle j_{y}\rangle_{\rm s}$. This is transformed as ${\cal Q}_{\rm c}=-{K}^{\rm(a)}\Delta T-({R^{\rm(b)}}/{2})\left(J_{y}-J_{y}^{*}\right)^{2}+({T_{\rm h}^{2}{S}^{\rm(a)\,2}_{yx}}/{2R^{\rm(b)}})(l_{y}/l_{x})^{2}$, where $J_{y}^{*}=-{T_{\rm h}{S}^{\rm(a)}_{yx}(l_{y}/l_{x})}/{R^{\rm(b)}}$. From this result we see that ${\cal Q}_{\rm c}$ is maximized for $J_{y}=J_{y}^{*}$, yielding the following maximum temperature difference
@@ III.4 Performance of transverse active cooler
where $a=2S_{yx}\langle j_{y}\rangle_{\rm s}/\kappa_{xx}$ and $b=\rho_{yy}\langle j_{y}\rangle_{\rm s}^{2}/\kappa_{xx}$. Note that this expression is also obtained by interchanging $T_{\rm h}$ and $T_{\rm c}$ in Eq. (74). Then, substituting this temperature distribution into Eq. (68), we obtain the rate of heat removal from the hot reservoir ${\cal Q}_{\rm h}=l_{y}l_{z}q_{x}(x=0)$ as ${\cal Q}_{\rm h}=K\Delta T-({R}/{2})J_{y}^{2}-T_{\rm c}S_{yx}(l_{y}/l_{x})J_{y}$, where $J_{y}=l_{x}l_{z}\langle j_{y}\rangle_{\rm s}$, and $K$ and $R$ are defined below Eq. (47). Completing the square with respect to $J_{y}$ in the above equation, we obtain ${\cal Q}_{\rm h}=K\Delta T-({R}/{2})\big(J_{y}-J_{y}^{*}\big)^{2}+({T_{\rm c}^{2}S_{yx}^{2}}/{2R})(l_{y}/l_{x})^{2}$, where $J_{y}^{*}=-T_{\rm c}S_{yx}(l_{y}/l_{x})^{2}/2R$. Then, by setting $J_{y}=J_{y}^{*}$, the maximum heat removal from the hot reservoir ${\cal Q}_{\rm h}^{*}$ is obtained as ${\cal Q}_{\rm h}^{*}=K\Delta T+({T_{\rm c}^{2}S_{yx}^{2}}/{2R})(l_{y}/l_{x})^{2}$. Now we introduce the heat current density $q_{\rm h}^{*}={\cal Q}^{*}_{\rm h}/(l_{y}l_{z})$, and rewrite the above equation as $q_{\rm h}^{*}=\kappa_{xx}\left(({\Delta T}/{l_{x}})\right)+({T_{\rm c}^{2}S_{yx}^{2}})/({2\rho_{yy}l_{x}})$. Then, defining $\widetilde{\nabla}_{x}T=\Delta T/l_{x}$ which is positive by definition, the above equation can be represented as
@@ IV.4 Transverse active cooler
Finally, we discuss the performance of the active cooler based on the off-diagonal Peltier effect. Let us first consider the case of isothermal condition in the transverse direction. In this case, as in the previous section, we need to determine the temperature distribution by solving Eq. (72). Then, we obtain the temperature distribution expressed by Eq. (82), where $a=0$ and $b=\rho_{yy}j_{y}^{2}/\kappa_{xx}$ should be noted. Then, substituting this temperature distribution into Eq. (68), we obtain the rate of heat removal from the hot reservoir ${\cal Q}_{\rm h}=l_{y}l_{z}q_{x}(x=0)$ as ${\cal Q}_{\rm h}=K\Delta T-({R}/{2})J_{y}^{2}+T_{\rm h}S_{yx}(l_{y}/l_{x})J_{y}$, where $J_{y}=l_{x}l_{z}j_{y}$. Completing the square with respect to $J_{y}$ in the above equation, we obtain ${\cal Q}_{\rm h}=K\Delta T-({R}/{2})\big(J_{y}-J_{y}^{*}\big)^{2}+({T_{\rm h}^{2}S_{yx}^{2}}/{2R})(l_{y}/l_{x})^{2}$, where $J_{y}^{*}=-{T_{\rm h}S_{yx}}(l_{y}/l_{x})/2R$. Then, by setting $J_{y}=J_{y}^{*}$, the maximum heat removal from the hot reservoir ${\cal Q}_{\rm h}^{*}$ is obtained as ${\cal Q}_{\rm h}^{*}=K\Delta T+({T_{\rm h}^{2}S_{yx}^{2}}/{2R})(l_{y}/l_{x})^{2}$. Now we define $q_{\rm h}^{*}={\cal Q}^{*}_{\rm h}/(l_{y}l_{z})$, and rewrite the above equation as $q_{\rm h}^{*}=\kappa_{xx}\left({\Delta T}/{l_{x}}\right)+({T_{\rm h}^{2}S_{yx}^{2}}/{2\rho_{yy}l_{x}})$. Then, with the definition of $\widetilde{\nabla}_{x}T=\Delta T/l_{x}$, the above equation can be transformed to Eq. (83), in which the effective thermal conductivity for the present case is given by
@@ V Measurement of transverse thermopower
The transverse thermoelectric effects can be investigated by measuring an electric field or voltage in a sample in the transverse direction ($y$ direction) while applying a uniform temperature gradient in the longitudinal direction ($x$ direction). In the measurement of the transverse thermoelectric effects due to structural symmetry breaking, which is independent of a magnetic field $H$ and/or magnetization $M$, the positions of electrodes for measuring the voltage must be carefully set up because asymmetry in the electrode positions cause the longitudinal thermopower to be superimposed on the transverse thermopower. It is easy to separate the transverse thermoelectric effects due to time-reversal symmetry breaking, which show an odd dependence on $H$ and/or $M$ in the $z$ direction, from the longitudinal thermopower. Here, the pure transverse thermopower can be extracted by measuring the $H$ dependence of the transverse voltage $V$ and calculating (V(+H)V(H))/2(V(+H)-V(-H))/2 with $V(+H)$ and V(H)V(-H) respectively being $V$ obtained when positive and negative magnetic fields are applied. In the case of the longitudinal thermoelectric effects, where the temperature gradient and electric field are parallel, the voltage per temperature difference and the electric field per temperature gradient are the same value. In contrast, in the case of the transverse thermoelectric effects, these are different values depending on the aspect ratio of the sample. Therefore, to discuss the performance of the transverse thermoelectric effects, the electric field per temperature gradient, i.e. the transverse thermopower, must be compared.
@@ V.1 Measurement configuration for Nernst effect in thin films
Figures 6(c) and 6(d) show an example of the experimental results of ANE measured in the PM and IM configurations, respectively, for a 50-nm-thick Co2MnGa thin film, epitaxially grown on a single-crystalline MgO (001) substrate [37]. In both the configurations, the observed transverse thermoelectric signals showed an odd dependence on $H$ and its magnitude saturated in the high-$H$ region, where the magnetization of the Co2MnGa film saturated. This is a typical behavior of ANE. In the IM configuration, a large ANE-induced thermopower was observed to appear even at zero field because the strong demagnetization field in the out-of-plane direction stabilized the remanent magnetization in the in-plane direction [Fig. 6(d)], a situation preferable for practical applications. In contrast, a tiny ANE-induced thermopower was observed at zero field in the PM configuration [Fig. 6(c)]. To generate the ANE-induced thermopower in the absence of an external magnetic field in the PM configuration, it is necessary to use a magnetic thin film having large perpendicular magnetic anisotropy [18, 55].
@@ V.2 Effect of thermal boundary conditions
We now experimentally demonstrate the influence of sample dimensions on the transverse thermopower using a bulk polycrystalline SmCo5 slab. The SmCo5 block, commercially available from Magfine Corporation, Japan, was cut into two rectangular slabs with dimensions along the $x$, $y$, and $z$ directions respectively being 10.0, 7.7, and 0.8 mm and 10.0, 2.1, and 0.8 mm using a diamond wire saw, where only the length along the $y$ direction, $d$, is significantly different. The slabs have an easy axis of magnetization along $z$ direction. The measurement configuration is illustrated in Fig. 7(c). The two SmCo5 slabs were bridged between two heat baths separated by 8.0 mm and firmly attached to the heat baths. To apply a temperature difference between the heat baths, chip heaters were attached to the hot side. The sample holder was fixed at the center of a superconducting magnet, where a magnetic field ranging from $\mu_{0}H=5~{\rm T}$ to $5~{\rm T}$ was applied along the $z$ direction. The $H$ dependence of the transverse voltage $V_{y}$ along the $y$ direction was recorded by attaching Cu wires at the ends of the slabs at various values of the temperature difference. To accurately measure the temperature gradient on the SmCo5 slabs, the top surface of the slabs was coated with insulating black ink with an emissivity of >0.94>0.94 and steady-state temperature images were captured using an infrared camera.
@@
Figure 7(d) shows the $H$ dependence of the transverse electric field $V_{y}/d$ for the SmCo5 slabs, recorded under a longitudinal temperature gradient of $|\nabla_{x}T|\sim 1\times 10^{3}~{\rm K/m}$. In both the slabs, the transverse electric field changed its sign with large hysteresis with respect to the $H$ reversal. The $H$ dependence of $V_{y}/d$ corresponds to the magnetization curve of the SmCo5 permanent magnet, indicating the appearance of ANE. To determine the magnitude of the $M$-dependent transverse thermopower for the SmCo5 slabs, the $M$-odd-dependent component of $V_{y}/d$ under the saturation state, $V_{\rm odd}/d$, was extracted for each $|\nabla_{x}T|$ value by extrapolating the signal in the high $H$ region to zero field. Here, the finite $H$-linear component was observed only for the slab with $d=2.1~{\rm mm}$, suggesting the presence of the Seebeck voltage between the SmCo5/Cu-wire contacts induced by the ordinary ($H$-dependent) THE in SmCo5 [Fig. 7(d)]. Although the SmCo5 slabs were cut from the same ingot, they exhibited different $V_{\rm odd}/d$ values [Fig. 7(e)]. The $M$-dependent transverse thermopower for the SmCo5 slab with $d=2.1~{\rm mm}$ ($d=7.7~{\rm mm}$) was found to be $5.1~\mu{\rm V/K}$ ($4.1~\mu{\rm V/K}$); the difference in the transverse thermopower is much larger than measurement errors. Recall the fact that the slab with $d=2.1~{\rm mm}$ ($d=7.7~{\rm mm}$) is closer to the adiabatic (isothermal) condition. Therefore, the transverse thermopower under the adiabatic condition is expected to be overestimated compared to that under the isothermal condition because of the same sign of $S_{xy}$ and the correction term induced by the Seebeck effect and anomalous ($M$-dependent) THE. This expectation is consistent with the observed difference in the $M$-dependent transverse thermopower. Here, we note that the correction term discussed theoretically in Sec. III is determined by the absolute Seebeck coefficient of a magnetic material (i.e. SmCo5), while the observed difference in the transverse thermopower in Fig. 7 depends on the relative Seebeck coefficient between SmCo5 and Cu. However, since the Seebeck coefficient of Cu ($2~{\rm\mu V/K}$) is much smaller than that of SmCo5 ($19~{\rm\mu V/K}$ [63]) at room temperature, the magnitude and sign of the correction term in the present setup is determined mainly by the transport properties of SmCo5. The experimental results shown here offer guidelines for determining the appropriate sample aspect ratio and measurement setup for transverse thermopower measurements, minimizing parasitic signals from the Seebeck effect and THE. We also note that considering typical sample dimensions, the estimation of $S_{\rm ANE}$ from the measurement of the anomalous Ettingshausen effect and the Onsager reciprocal relation reduces the influence of the correction term [29].